4. The money base is $300, the reserve requirement is 5%, banks hold no excess reserves and the public holds no currency.

a. What is M2?

b. If the velocity of M2 is 4 (and this is independent of output and the price level), write a function representing aggregate demand for this economy.

c. If the long-run equilibrium level of output is 400, what is the equilibrium price level?

d. Show what will happen in both the short run and the long run if the central bank sells $50 worth of bonds. What will happen to the unemployment rate in the short run and the long run? What will the price level become in the long run?
Answer #1

A) M2 = MB* m

m = multiplier = 1/RRR =1/.05 = 20

M2 = 300*20 = 6,000

B) MV = PY , from Quantity theory of money

AD : PY = 6000*4 = 24,000

So P = 24,000/Y

c) Y = 400

p* = 24,000/400 = 60

D) CB sells 50 $ bonds , so it withdraws money from the economy,

Thus it's a contractionary policy, thus unemployment will rise in both short & long run

Thus %∆ in P = %∆ in Money supply